OpenStudy (anonymous):

A wallet containing £40, has 3 times as many £1 as £5. find the number of each kind.

OpenStudy (lgbasallote):

what is 3 times as many 1 euro and 5 euro?

OpenStudy (anonymous):

re worded the question, I was typing in anger.

Parth (parthkohli):

3(1x) + 5x = 40 3x + 5x = 40 8x = 40 x = 40/8 = 5 So, There are 15 one euro notes and 5 five euro notes.

Parth (parthkohli):

Do you get it?

OpenStudy (lgbasallote):

ohh i see

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

actually, I cant see how (3(1x) and 5x) represent that there are 3 times as many 1s then 5s?

Parth (parthkohli):

Exactly! You got it now!

OpenStudy (anonymous):

I do not understand

OpenStudy (henryblah):

x is the amount of £5 notes.

Parth (parthkohli):

Okay, let me explain. THERE ARE 3 TIMES 1 EURO NOTES THAN THE 5 EURO NOTES. THE TOTAL IS 40 EUROS. 1(3x) + 5(x) = 40 x represents the number of notes. From the above equation, we see that there are 1 euro notes 3 times the 5 euro notes. Now you know it :D

OpenStudy (anonymous):

That all good, When I see 3x + 5x but I do not equate that to one number being 3 times another.

OpenStudy (anonymous):

that is really what is baffling me.

Parth (parthkohli):

We can say that 5x and 3x are like terms. The coefficient of the numerical value is same, so you can add. Similarly, you can add: \(2x^{2} + 5x^{2} \) \(4x^{2}y^{3} + 6x^{2}y^{3} \) etc. etc.

OpenStudy (henryblah):

Ok. If you have the same number of £1 and £5, let's say x. What is the total amount of money? Would you agree that is 1x+5x or (1+5)x? Now apply that to here, but instead you have 3 times the £1. So total money is 1*3x+5x. Then you put that equal to £40 to find x.

OpenStudy (anonymous):

@ henryblah, that makes a little more sense.